# Reduced Data-Driven Turbulence Closure for Capturing Long-Term Statistics

### v1

@@ Abstract
We introduce a simple, stochastic, a-posteriori, turbulence closure model based on a reduced subgrid scale term. This subgrid scale term is tailor-made to capture the statistics of a small set of spatially-integrate quantities of interest (QoIs), with only one unresolved scalar time series per QoI. In contrast to other data-driven surrogates the dimension of the “learning problem” is reduced from an evolving field to one scalar time series per QoI. We use an a-posteriori, nudging approach to find the distribution of the scalar series over time. This approach has the advantage of taking the interaction between the solver and the surrogate into account. A stochastic surrogate parametrization is obtained by random sampling from the found distribution for the scalar time series. Compared to an a-priori trained convolutional neural network, evaluating the new method is computationally much cheaper and gives similar long-term statistics.
@@ 1 Introduction
The forced barotropic vorticity equation is used as test bed, which is widely used to test (data-driven) SGS models [33, 35, 12, 13, 32] and can be seen as one of the simplest models in large-scale atmospheric and oceanic modelling when temperature variations are not significant.
@@ 2.1 Spectral discretization
where $\hat{\omega}_{\boldsymbol{k}}$ is the Fourier coefficient for the wave number vector $\boldsymbol{k}=(k_{1},k_{2})^{T}$ and $N$ is the number of grid points in each direction of the square grid. Time integration is performed using the standard fourth-order Runge-Kutta scheme. The non-linear term $J$ is computed with a pseudospectral technique, using the “3/2 rule” to remove aliasing errors, see B.
@@ 2.2 Multiscale decomposition
We apply a spectral filter to a high-fidelity (HF) reference simulation, to obtain a low-fidelity (LF) model with a lower spatial resolution. The HF model resolves $257\times 257$ Fourier modes ($N=257$). A time step of $\Delta t=0.01$ days is used. Our LF simulations resolve $65\times 65$ Fourier modes instead ($N_{LF}=65$), and uses a 10 times larger time step: $\Delta t=0.1$. The spectral filter simply removes all wave numbers from the HF simulations which are not present in the LF model:
@@
Thus, $C=\lfloor\frac{N_{LF}}{2}\rfloor=32$ is the cutoff frequency, and $\hat{R}_{C}$ is a square filter, see Figure 1(a). In the following, we will write $\bar{x}=\mathcal{F}^{-1}\left(\hat{R}_{C}\hat{x}_{\bf k}\right)$, such that $\bar{x}$ quantity in the physical domain, filtered by $\hat{R}_{C}$ in the spectral domain. Applying this filter to (1) gives
@@ 3.2 CNN
We use the CNN architecture tuned in [12], and applied to essentially the same governing equations as studied here. It contains 10 convolutional layers with convolutional depth 64 and filter size $5\times 5$. ReLU activation functions are used, except for the linear output layer. The only difference is that our layers use periodic padding, in line with our periodic boundary conditions, whereas [12] used zero padding. The CNN is trained to approximate the mapping
@@ 5.1.2 Convergence of distributions
To get an intuition of what can be considered a small KS-distance, we compare the marginals of our QoIs over the first 10000 days of a 20000-day HF simulation to the distribution over the full 20000 days, see Table 1.
@@ 5.1.3 Tau-orthogonal prediction results
    Multivariate Gaussian sampling: we sample form a multivariate Gaussian distribution (MVG) which was fitted to the joint distribution of the corrections $\Delta Q_{i}$.
@@ 5.4 Comparison of subgrid scale surrogates
Figure 12:
@@ 5.4.1 Computational cost
The CNN subgrid parameterization leads to serious overhead. In fact, one LF time step with CNN is slower than one HF time step, such that all the speedup can be attributed to the coarser time step. Normally, making predictions in batches improves efficiency of neural networks. However, due to the recurrent nature of predicting the subgrid term this is not possible.
@@
Before the CNN and TO surrogates can be used, they need to be trained on reference data. Table 3 reports the training time for a CNN and the TO parameterization. Note that training the TO method involves running an LF simulation that tracks the HF QoI data in the training domain. The TO method needs a longer reference run than the CNN, which leads to extra costs. Taking all this together the total runtime for finding long term distributions is roughly the same with the TO method and the CNN parametrization. Note that this is problem specific and heavily depends on how much training data both methods need.
@@ 6 Conclusion
A downside of the tau-orthogonal method is that (by construction) it cannot give guarantees about the long-term statistics of quantities outside the predefined set of QoIs. One has to choose beforehand what must be learned from the data. We also observed that with the CNN SGS model, the vorticity fields retain correlation with the high-fidelity reference fields for a longer period. Mainly because the local spatial correlation of the vorticity fields is better captured when using the CNN SGS model. We expect that this is due to the choice of spatio-temporal basis functions in the tau-orthogonal method, which is the subject of future research.
Another potential research direction is the application of the tau-orthogonal method to three-dimensional (flow) problems. Here the computational costs and memory requirements for a CNN parameterization increase tremendously due to the cubic scaling of the problem size. On the other hand, the dimension of the learning problem for the tau-orthogonal method is independent of the spatial dimension of the differential equation under consideration.
@@
The implementation of the methods which are presented in this paper is publicly available at [15]. This code repository includes a notebook, which takes the user step by step through the experiments in the paper.

### v2

@@ Abstract
We introduce a simple, stochastic, a-posteriori, turbulence closure model based on a reduced subgrid scale term. This subgrid scale term is tailor-made to capture the statistics of a small set of spatially-integrated quantities of interest (QoIs), with only one unresolved scalar time series per QoI. In contrast to other data-driven surrogates the dimension of the “learning problem” is reduced from an evolving field to one scalar time series per QoI. We use an a-posteriori, nudging approach to find the distribution of the scalar series over time. This approach has the advantage of taking the interaction between the solver and the surrogate into account. A stochastic surrogate parametrization is obtained by random sampling from the found distribution for the scalar time series. We compare the new method to an a-priori trained convolutional neural network on two-dimensional forced turbulence. Evaluating the new method is computationally much cheaper and gives similar long-term statistics.
@@ 1 Introduction
The forced barotropic vorticity equation is used as test bed, which is widely used to evaluate (data-driven) SGS models [32, 34, 12, 13, 31] and can be seen as one of the simplest models in large-scale atmospheric and oceanic modelling when temperature variations are not significant. While this application thus has some relevance, we reiterate that the main purpose of this article is to evaluate the predictive capability of tau-orthogonal method on (spatially-integrated) QoIs, using an as small-and-simple data-driven surrogate as possible, especially compared to a CNN with a multitude of tunable parameters.
@@ 2.1 Spectral discretization
where $\hat{\omega}_{\boldsymbol{k}}$ is the Fourier coefficient for the wave number vector $\boldsymbol{k}=(k_{1},k_{2})^{T}$ and $N$ is the number of grid points in each direction of the square grid. Time integration is performed using the standard fourth-order Runge-Kutta scheme. The non-linear term $J$ is computed with a pseudospectral technique, using the “3/2 rule” to remove aliasing errors, see B. The discrete system is solved using double precision arithmetic.
@@ 2.2 Multiscale decomposition
We aim to learn a subgrid scale term to improve the accuracy of a computationally cheap low-fidelity (LF) model, using data from a more accurate high-fidelity (HF) model. The LF model uses $65\times 65$ ($N=65$) Fourier modes, such that it resolves only the largest scales of the problem. We use a HF model with $257\times 257$ ($N=257$) Fourier modes as ground truth, following [34] and [4]. With this setup the HF model resolves a significantly larger part of the turbulence spectrum than the LF model. Note that it is not feasible to run a fully resolved direct numerical simulation of the problem due to the high Reynolds number, $Re\sim O(10^{6})$. The HF model uses a time step of $\Delta t=0.01$ days. Our LF simulations use a 10 times larger time step: $\Delta t=0.1$. Low fidelity fields are obtained with a spectral filter, simply removing all wave numbers from the HF simulation which are not present in the LF model:
@@
Thus, $C=\lfloor\frac{N_{LF}}{2}\rfloor=32$ is the cutoff frequency, and $\hat{R}_{C}$ is a square filter, see Figure 1(a). In the following, we will write $\bar{x}=\mathcal{F}^{-1}\left(\hat{R}_{C}\hat{x}_{\bf k}\right)$, such that $\bar{x}$ is a quantity in the physical domain, filtered by $\hat{R}_{C}$ in the spectral domain. Applying this filter to (1) gives
@@ 3.2 CNN
We use the CNN architecture tuned in [12], and applied to essentially the same governing equations as studied here. It contains 10 convolutional layers with convolutional depth 64 and filter size $5\times 5$. ReLU activation functions are used, except for the linear output layer. It is evaluated using single precision arithmetic, as is common practice in machine learning applications. The only difference is that our layers use periodic padding, in line with our periodic boundary conditions, whereas [12] used zero padding. The CNN is trained to approximate the mapping
@@ 5.1.2 Convergence of distributions
Instead of using a pointwise measure as (41), it might be appropriate to use a similarity measure that integrates over the support of the two pdfs, e.g. the Hellinger or Wasserstein distance. However, to be able to compare the distances obtained from different QoIs binning or normalization choices have to be made, because the distributions of the QoI live on different scales. We opted to retain the simpler KS statistic, avoiding these choices. We found our results to be robust under a change to the Hellinger distance (where we binned based on the spread of the reference distribution). To get an intuition of what can be considered a small KS-distance, we compare the marginals of our QoIs over the first 10000 days of a 20000-day HF simulation to the distribution over the full 20000 days, see Table 1.
@@ 5.1.3 Tau-orthogonal prediction results
    Multivariate Gaussian sampling: we sample from a multivariate Gaussian distribution (MVG) which was fitted to the joint distribution of the corrections $\Delta Q_{i}$.
@@ 5.4 Comparison of subgrid scale surrogates
Figure 12: Long-term probability-density functions of the 4 QoI.
@@ 5.4.1 Computational cost
The CNN subgrid parameterization leads to serious overhead. In fact, one LF time step with CNN is slower than one HF time step, such that all the speedup can be attributed to the coarser time step. Normally, making predictions in batches improves efficiency of neural networks. However, the input of the CNN in the subgrid parametrization depends on the output of the CNN in the previous time step. Due to this recurrent nature of predicting the subgrid term, it is not possible to predict the SGS term for multiple time steps in one batch.
@@
Before the CNN and TO surrogates can be used, they need to be trained on reference data. Table 3 reports the training time for a CNN and the TO parameterization. Note that training the TO method involves running an LF simulation that tracks the HF QoI data in the training domain. The TO method needs a longer reference run than the CNN, which leads to extra costs. Taking all this together, the total runtime for finding long term distributions is roughly the same with the TO method and the CNN parametrization. Note that this is problem specific and heavily depends on how much training data both methods need. Also the efficiency of the code implementation and the compute architecture play an important role. Our base solver is an in-house Python code, certainly slower than (more optimized) similar codes written in other languages. Moreover, newer GPUs tend to be optimized for deep learning, leading to faster training and evaluation of the CNNs. The evaluation time of the base solver and the tau-orthogonal method (which also run on the GPU), would likely not profit at the same rate from these optimizations, which would alter the reported relative performance numbers.
@@ 6 Conclusion
A downside of the tau-orthogonal method is that (by construction) it cannot give guarantees about the long-term statistics of quantities outside the predefined set of QoIs. One has to choose beforehand what must be learned from the data. We also observed that with the CNN SGS model, the vorticity fields retain correlation with the high-fidelity reference fields for a longer period. This is mainly because the local spatial correlation of the vorticity fields is better captured when using the CNN SGS model. We expect that this is due to the choice of spatio-temporal basis functions in the tau-orthogonal method, which is the subject of future research.
Thus far, the tau-orthogonal method has only been applied to the 2D barotropic Navier-Stokes equations. While this has relevance for geophysical flow simulations, an essential research direction is the application to 3D flow problems. Such a realistic flow case is important to truly evaluate the promise of the data-driven method. Here, the computational costs and memory requirements for the base solver increase tremendously, due to the cubic scaling of the problem size. This means that the relative decrease in computational cost from coarse graining also is significantly larger than in the two-dimensional case, i.e. scaling cubically instead of quadratically. CNN parameterizations for such problems rely on three-dimensional kernels, drastically increasing the network size and training cost compared to their two-dimensional counter parts. On the other hand, in the tau-orthogonal method the dimension of the learning problem is independent of the spatial dimension of the differential equation under consideration. Performance wise, the only significant extra costs are incurred by taking inner products of three-dimensional vector fields. Whether the tau-orthogonal method is able to accurately capture quantities of interest in a large 3D simulation is the subject of ongoing research.
@@
The implementation of the methods which are presented in this paper is publicly available at [hoekstra_2024]. This code repository includes a notebook, which takes the user step by step through the experiments in the paper.